OpenStudy (countrygirl1431):

Evaluate [3(5 + 6) + 2] ÷ 7 3 5

OpenStudy (sweetburger):

PEMDAS

OpenStudy (countrygirl1431):

?????

OpenStudy (jjwalton04):

3(5+6)+2/7 = (3)(11)+2/7 = 33+2/7 = 35/7 =5

OpenStudy (amilapsn):



OpenStudy (jjwalton04):

lol kinda ddefeats the purpose though